### v5

# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states. These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling. We demonstrate our approach on two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms. Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study. Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in less than 15 minutes on a standard GPU rather than weeks on specialized hardware.
1 Introduction
Characterizing molecular conformational changes, or transitions, is essential for understanding biological processes, from enzyme catalysis to protein folding. Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes. This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935). As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive.
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as splines or as a discrete series of molecular configurations connected by artificial spring forces to ensure continuity. Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for small-scale systems, it has not yet benefited from modern optimization approaches that could address these challenges.
Recent advances in machine learning methodology have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical force field models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives beyond forces, which are needed for neural path optimization.
Numerous approaches have emerged over the past decades to tackle the problem of sampling transitions. Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics to begin with. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Lee et al., 2025; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
Notably, Du et al. (2024) uses a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024). Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, a class of methods related to string methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics to optimize the neural path. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a force field to generate plausible transition paths using the discretized Onsager-Machlup action directly as a loss function. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to much larger biomolecular systems with explicit solvent, essential for accurately modeling such systems, while maintaining fast training times.
Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy-path finding problem as a continuous optimization framework. Our approach transforms the Onsager-Machlup functional, which naturally models the probability of paths in stochastic systems, into a remarkably simple and intuitive loss function that requires only energy evaluations along the neural network parametrized path. In contrast to existing methods that use complex composite losses involving energies, forces, and path gradients, our streamlined formulation eliminates the need for these additional computations. This simplification, while faster, also sidesteps known issues in PINN optimization if the loss contains differential operators (Rathore et al., 2024). This PINN framework efficiently captures physically meaningful transition mechanisms for complex biomolecular systems, completing optimization in minutes on standard hardware.
2 Theory and Methods
2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To derive our loss function, which when minimized yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods. However, we further simplify the approach after reaching their geometric formulation. A complete derivation with all mathematical details is provided in Appendix A.1. We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
represents the potential energy function, is Gaussian white noise modeling thermal fluctuations, and physical constants are absorbed into the nondimensionalization. To derive the path probability, we discretize the time interval into steps of duration , approximating the dynamics as:
where are independent standard Gaussian random variables. The conditional probability of transitioning from to follows a Gaussian distribution:
Taking the product of these conditional probabilities over all time steps and then taking the limit as yields the path probability:
The exponent corresponds to the Onsager-Machlup action functional Onsager & Machlup (1953):
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. Notably, there are different derivations of the action that yield an additional higher-order derivative term related to trajectory entropy, whose necessity is conditional on the task (Adib, 2008). Since we are primarily interested in MEPs, we neglect this term. However, the resultant MEPs still empirically align well with the free energy landscape. To further simplify, we expand the integrand of the action functional:
For paths with fixed endpoints, the cross-term integrates to a constant difference in potential energy:
Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization. We can therefore focus on minimizing the remaining terms. Applying the Cauchy-Schwarz inequality to the remaining terms:
This upper bound becomes an equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape. Crucially, we allow the path to optimize its shape and length in our formulation. This freedom enables the path to adjust to follow the energy gradient (or its opposite), making the above upper bound tight at the optimum. When we abandon fixed-time parameterization and allow the path to stretch as needed to align with energy gradients, we can transform the time integral into a purely geometric one by defining the arc-length element along the spatial curve :
While the referenced works (Vanden-Eijnden & Heymann, 2008; Olender & Elber, 1997) derive and use this geometric action, we further simplify it. This geometric integral represents a natural endpoint for theoretical development for traditional chain-of-states and string methods. In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string using forces conceptually straightforward.
However, we want a loss function that acts directly on the neural network parameters in our neural network formulation. While we could parameterize the path using a neural network and then compute forces along the path before backpropagating to our parameters, this would introduce additional gradient computations, which have been noted to lead to optimization challenges in PINNs (Rathore et al., 2024). We avoid this issue by simplifying to an energy-based loss.
We represent the transition path as a continuous neural mapping from a progress coordinate to the system’s Cartesian coordinates, such that and . We then discretize the path into segments:
where and .
For a path segment, applying Taylor’s theorem, we can express the energy difference as:
Where is the angle between the gradient and displacement vectors. For MEPs, the displacement aligns with or against the gradient direction () giving:
This gives us a surrogate objective with the same minimum:
Using the triangle inequality on the absolute energy differences, we can establish that minimizing the simple sum of energies along the path provides an upper bound on this action:
Where is the lowest energy along the path. Since the latter terms are constants during optimization, this leads to our final loss function. To implement this in practice, we sample B points along a parametrized path and directly minimize the sum of energies at these points, yielding:
where are uniformly or randomly sampled points in . By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
2.1.1 Neural Representation of MEPs
We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms, as illustrated in Figure 1 (left):
Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the neural network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the neural network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
2.1.2 Path Optimization
We minimize using the Adam optimizer (Kingma & Ba, 2017):
Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary; the inherent continuity of , combined with smooth blending functions, ensures path continuity.
2.2 Established Approaches to MEP Finding
Classical approaches to MEP discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation. This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization. The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as by E & Vanden-Eijnden (2010). These established techniques illustrate both foundational concepts and limitations that our approach addresses.
2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized using splines mapping to the systems Cartesian coordinates that connects these states. The MEP satisfies:
where is the component of the force perpendicular to the unit tangent vector along the string. This orthogonality condition has an intuitive interpretation: at each point along the MEP, the forces acting on the system are entirely parallel to the path itself, with no perpendicular components trying to push the system away from the path. This represents a ridge or valley in the energy landscape—the system naturally follows the path without being forced sideways.
The simplest dynamics for evolving a curve toward this MEP is given by:
Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all . In practice, the string is discretized into images and evolved through a two-step procedure that requires explicit reparameterization steps:
1. Evolution step:
2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition .
2.2.2 Chain-of-States Methods
An alternative perspective views MEP finding as optimization of a chain-of-states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008). Given a sequence with and , the objective function is defined as:
where and is a spring constant. The spring term enforces the continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost.
3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025). In Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, neural network architectures, and regularization approaches.
3.1 AIB9 Transition Path Discovery
We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods. This system effectively bridges abstract models and complex proteins: it is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes; its energy landscape is well-characterized; and its transitions are easily visualized through dihedral angle projections. Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields. We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space. The method identifies multiple physically plausible pathways between the states, which are also observed with long MD simulations. For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions. Notably, the discovered MEPs align well with the free energy landscape despite our choice of action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.
4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. BPTI is 58 residues long (892 atoms) and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent larger system to test the method. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
We generate multiple transition MEPs connecting two BPTI states from the D.E. Shaw simulation. The intermediate snapshots from the D.E. Shaw simulation serve as validation points; they were not used as inputs but represent physically realized conformations from the whole trajectory. To analyze this path, we compute two collective variables matching those used in the original D.E. Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD, root mean square displacement, of residues 4-54 after mean-centering. Remarkably, as shown in Figure 3, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive MD simulation.
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. This increases the number of atoms from 892 to 8,300. To handle water molecules during transitions, we pair proximal water molecules between start and end states using the Hungarian algorithm (Kuhn, 1955). We then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996). Each MEP was computed in approximately 15 minutes on a single A6000 GPU, requiring only force field evaluations compared to the D.E. Shaw simulation, which conducted approximately 412 billion force field evaluations to generate the 1.03 millisecond trajectory capturing the same conformational changes. This represents a six orders of magnitude reduction in computational effort.
While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway, which is often the primary mechanistic insight sought by MD practitioners. For most scientifically interesting systems—typically large biomolecular complexes—equilibrium sampling of the complete conformational space or transition path ensemble remains computationally prohibitive, making identification of dominant pathways particularly valuable. It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space. Appendix B.2.7 provides a detailed computational performance comparison.
5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach, outlined in B.2.5, largely mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies. Our approach is also constrained to systems with suitable differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying MEPs rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system. However, as noted in the BPTI section 4, this mechanistic insight is often a key objective for MD practitioners studying complex biomolecular transitions.
6 Conclusion
We demonstrated that reformulating MEP discovery as a continuous neural optimization problem enables efficient discovery of molecular transition mechanisms with unprecedented computational efficiency. Our approach transforms the complex Onsager-Machlup action functional into a straightforward energy-based loss function that eliminates the need for force and path gradient computations required by related methods, which are known to induce optimization challenges in PINNs. This work provides a practical tool for molecular dynamics practitioners that captures essential mechanistic insights without requiring specialized hardware or extensive simulation time.
While our current implementation shows promising results, several concrete directions for improvement remain. Future work should explore neural architectures that incorporate domain-specific structural priors beyond simple MLPs with sinusoidal/Fourier embeddings. Related work by Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path that might further aid convergence for paths that constitute dramatic rearrangements. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from conformations that already respect molecular geometry. Extending the method to handle larger multi-domain proteins is also a frontier of great scientific interest, which might bring new challenges. Future work could also explore ways to estimate free energies and reaction rates from our discovered paths.
7 Broader Impacts
Our research democratizes access to the study of larger molecular systems by enabling transition path studies on standard hardware. This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption. We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could be misused in the future, and we encourage responsible use with appropriate experimental validation.
Appendix A Technical Appendices and Supplementary Material
A.1 Extended Derivation of the energy-based Loss Function
In this appendix, we provide a comprehensive derivation of our energy-based loss function for discovering minimum-energy paths (MEPs) between molecular configurations. The derivation proceeds in four main steps: (1) establishing the path probability density from overdamped Langevin dynamics, (2) reformulating this probability in terms of the Onsager-Machlup action functional, (3) simplifying this functional to a geometric line integral, and (4) developing a discretized approximation suitable for neural network optimization.
Our goal is to find the most probable transition path between two stable molecular configurations and . This path corresponds to the minimum-energy pathway through the potential energy landscape and represents the most likely mechanism of conformational change.
Assumptions and Approximations
The derivation relies on several key assumptions:
The molecular system evolves according to overdamped Langevin dynamics, appropriate for high-friction biomolecular environments
The path endpoints and are fixed and correspond to stable states
The potential energy function is differentiable everywhere along the path
We can represent the path as a continuous, parameterized curve in configuration space with sufficient flexibility
The path’s parameterization allows for effective length changes, permitting the path to adjust its stretching as needed to follow energy gradients optimally
We also make several important approximations during the derivation:
We use a first-order Taylor expansion to relate energy differences to gradient magnitudes
We introduce stochastic sampling to estimate the continuous energy integral
For a MEP, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
We use the triangle inequality to establish an upper bound on the sum of absolute energy differences
While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached.
A.1.1 From the Onsager-Machlup Action to a geometric Line Integral
We begin with a molecular system evolving under overdamped Langevin dynamics, a standard model for high-friction molecular motion:
where represents the molecular configuration, is the potential energy function, and is a Gaussian white noise with .
To derive the path probability, we discretize the time interval into steps of duration , approximating the dynamics as:
where are independent standard Gaussian random variables with zero mean and unit variance.
The conditional probability of transitioning from to follows a Gaussian distribution:
where is the dimensionality of the system. The probability of the entire discretized path is given by the product of these conditional probabilities:
Taking the limit as and with fixed, this sum approaches a path integral:
Thus, the path probability can be written as:
The exponent term is recognized as the Onsager-Machlup action functional:
This action functional penalizes paths that deviate from following the force . Maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional.
Expanding the integrand:
For paths with fixed endpoints and , the cross-term integrates to a constant difference in potential energy:
We can define a modified action functional that, when minimized, is equivalent to minimizing the original Onsager-Machlup functional for fixed endpoints:
By the Cauchy-Schwarz inequality:
with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape. This is precisely the defining characteristic of the MEP - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
When this equality holds, our action simplifies to:
The term represents the infinitesimal arc length element along the path, allowing us to rewrite the action as a line integral:
where is the path in configuration space. This geometric form is independent of the path’s parameterization and depends solely on the potential energy landscape.
A.1.2 Discretization and Connection to Energy Differences
To compute this geometric action numerically, we introduce a progress coordinate and a mapping such that and , where represents the network parameters. The relationship between the progress coordinate and arc length is:
We discretize the path with points, setting and for . The discretized geometric action becomes:
where .
Using Taylor’s theorem, we can relate energy differences to gradient magnitudes:
where is the remainder of the order and is the angle between the gradient and displacement vectors.
For a path following the minimum-energy trajectory, the gradient aligns with or against the path direction ( as , depending on whether the path is ascending or descending the energy landscape), giving at the minimum of the optimization:
This yields an approximation of the geometric action:
This forms a surrogate objective with the same minimum as our original action.
A.1.3 Establishing the energy-based Loss Function
We need to bound the sum of absolute energy differences to establish a direct connection to an energy-based loss function. Using the triangle inequality, for any two points and any constant :
Choosing to be the minimum energy along the entire path and summing over all segments:
We can rewrite the first sum as and the second as . Combining these terms:
Since , , and are constants during optimization, minimizing is equivalent to minimizing the upper bound on the geometric action. This provides the theoretical justification for our energy-based loss function.
A.1.4 Neural Network Implementation
For practical optimization, we parametrize the path using a neural network with parameters . To implement our approach, we sample points along the path and minimize the energy at these sampled locations. This represents a transition from the theoretical derivation using discrete segments to a practical implementation using sampled points for more efficient optimization. We employ stochastic sampling of the progress coordinate to construct our loss function:
where are uniformly sampled progress coordinates and is the batch size (the number of points sampled along the path during each optimization step).
This stochastic approach provides an unbiased estimator of the expected energy along the path:
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the MEP predicted by the Onsager-Machlup framework. The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
A.2 Connection between Doob’s Lagrangian and MEP Optimization
We elaborate on the theoretical connections between our energy-based optimization approach and the framework of Doob’s h-transform for sampling transition paths (Du et al., 2024). This comparison is particularly relevant as both methods employ neural networks to represent molecular state transitions—either as a minimum-energy path in our case or the full transition path ensemble in theirs. Notably, the parametrization of the mean trajectory in Du et al. resembles our MEP representation, differing primarily in our use of blending functions and spatial embeddings for the neural network component.
Doob’s Lagrangian formulation considers the following action functional:
This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape. Intuitively, it measures the amount of "effort"/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
In this formulation, is the probability density at time given boundary conditions, parameterized as a Gaussian , and is the control vector field satisfying , with representing the reference drift for overdamped dynamics.
This action functional describes a Wasserstein Lagrangian flow as formalized in (Neklyudov et al., 2023) and (Neklyudov et al., 2024). In this framework, one can learn stochastic dynamics from samples, thereby learning a process that defines a time-dependent density evolving from initial to final states.
The key difference with our PINN-MEP approach is that we minimize the action of a single deterministic path rather than a full distribution of paths. Mathematically, this corresponds to working with a single trajectory through configuration space rather than an ensemble of possible trajectories. This simplification is reflected in our loss function, which contains one fewer expectation than the full Doob’s Lagrangian.
A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work Lee et al. (2025). The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
Method | States | # Evaluations | Max Energy | MinMax Energy |
MCMC variable length | CV | 21.02M | 740.70 695.79 | 52.37 |
MCMC* | CV | 1.29B | 288.46 128.31 | 60.52 |
MCMC variable length | relaxed | 187.54M | 412.65 334.70 | 26.97 |
MCMC | relaxed | 10B | N/A | N/A |
MCMC variable length | exact | 10B | N/A | N/A |
MCMC | exact | 10B | N/A | N/A |
Doob’s Cartesian | exact | 38.40M | 726.40 0.07 | 726.18 |
Doob’s Cartesian, 2 Mixtures | exact | 51.20M | 709.38 162.37 | 513.72 |
Doob’s Cartesian, 5 Mixtures | exact | 51.20M | 541.26 278.20 | 247.96 |
Doob’s Internal | exact | 38.40M | -14.62 0.02 | -14.67 |
Doob’s Internal, 2 Mixtures | exact | 51.20M | -15.38 0.14 | -15.54 |
Doob’s Internal, 5 Mixtures | exact | 51.20M | -15.50 0.31 | -15.95 |
Doob’s Seq2Seq | exact | - | 1505.6 0.45 | 245.05 0.02 |
w/ Temperature Annealing | exact | - | 1583.18 0.3 | 592.13 0.26 |
w/ Interatomic Interpolation | exact | - | 1601.57 0.53 | 3.46 0.03 |
MaxLL | exact | - | 1532.45 | 615 |
w/ Temperature Annealing | exact | - | 1599.03 | 233 |
w/ Interatomic Interpolation | exact | - | 1545.32 | 619 |
PINN-MEP | exact | 15700 | -67.22 3.55 | -70.78 |
PINN-MEP requires approximately 3261× fewer energy function evaluations (15.7K vs. 51.20M) compared to the Doob’s Lagrangian approach while achieving a 4.44 times lower maximum energy barrier. The lower number of required evaluations is likely due to the computation of the loss, which requires sampling their path over both , as we do as well, but also the width of the path. Thus, the tradeoff of sacrificing full ensemble characterisation is expected to come with computational efficiency.
Appendix B Comparison with other related Methods
While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/MEP generation is also informative. Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
Method | # Evaluations | Max Energy (eV) |
NEB | 8063 | 11.0014 |
INR | 3046 | 0.4070 |
Method | CVs | # Evaluations | Runtime / Path |
MCMC (Two-Way Shooting) | No | 1B | 100 hours |
Metadynamics | Yes | 1M | 10 hours |
OM Opt. (Diffusion Model) | No | 10K | 50 min |
Table 2 shows results from Ramakrishnan et al. (2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for MEP discovery. Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields. Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity), but needs fewer force field evaluations than our method. It is important to note that the "Max Energy" reported in Table 2 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging. The difference in force fields used further complicates any direct energy value comparisons between methods.
In contrast, Table 3 presents results from Raja et al. (2025), who take a fundamentally different approach by re-purposing pre-trained generative models to perform sampling of transition paths. A key advantage of their method is that the generative model provides significantly better initial guesses for transition paths based on learned molecular distributions. Their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function, which is more computationally intense due to the high number of parameters of the pretrained model as compared to a physics-derived force field.
These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
B.1 Ablation Studies
This section presents a series of ablation studies examining the impact of various components of our PINN-MEP approach. These experiments help validate our design choices and provide insights into the importance of different architectural and training elements. We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method. For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the energy, and the mean number of iterations to reach the lowest maximum energy. We run each setting in the ablations 15 times for a maximum of 25,000 iterations. We use a batch size of 1, where each batch contains 32 values of the progress coordinate for most cases. The exception is when we evaluate the chain-of-states-like regularization, which minimizes the difference between network outputs for adjacent values of ; in this case, the effective batch size is 32, since there are no interactions between different network outputs otherwise.
Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies. This can limit the ability to compare the performance of different settings. However, we find that the results are still directionally informative.
B.1.1 General Settings for Ablation Studies
All ablation studies are conducted using the following base configuration, with each study varying only specific parameters while keeping all others fixed:
Molecular System Parameters:
Molecular system: AIB9 peptide with 129 atoms
Force field: AMBER ff15ipq-m for protein mimetics
Cutoff method: No cutoff applied
Simulation box size: 10.0 nm
Neighbor list cutoff distance: 4.0 nm
Training Configuration:
Number of iterations: 25,000 per run
Number of independent runs: 15 per parameter setting
Evaluation frequency: Every 100 training steps
Optimizer: Adam with optax default parameters
Learning rate:
Path Sampling Parameters:
Batch size: 1
Frames per batch: 32 points sampled along the progress coordinate
Spring constant: (no chain-of-states regularization)
Interpolation method: Cubic Hermite for blending functions
Network Architecture (Base Configuration):
Network depth: 6 hidden layers
Width factor: 4 (network width scales as number of atoms)
Embedding dimension: Automatically set to
Time embedding: Sinusoidal with maximum period of
Activation function: ReLU throughout the network
Weight initialization: Glorot uniform for hidden layers, zero initialization for final layer
Bias initialization: Default for hidden layers, zero initialization for final layer
B.1.2 Embedding Strategy
We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress coordinate. Table 4 shows that linear embedding (None) performed best, achieving the lowest mean energy, though sinusoidal embeddings with period also achieved competitive results. We find these results at odds with the need for sinusoidal embeddings in the larger BPTI system. Perhaps the AIB9 system, which is characterized by one large general conformational rearrangement primarily driven by main-chain dihedral angles, does not need the high frequency features that sinusoidal embeddings provide, but which might be essential for other systems, due to the well-known bias of MLPs toward modeling low frequencies in data.
Embedding | Max Period | MeanMax Energy () | MinMax Energy | Mean Iterations |
Linear (None) | – | 591.35 963.72 | -337.36 | 200873 |
Sinusoidal | 1326.13 2339.35 | -337.36 | 210907 | |
Sinusoidal | 1528.18 2429.05 | -129.90 | 226290 | |
Sinusoidal | 1888.66 3604.47 | -337.36 | 187493 | |
Sinusoidal | 2453.53 3507.47 | -337.36 | 217053 |
B.1.3 Spring Regularization
We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5. Interestingly, we found that small amounts of spring regularization ( or ) provided the best results, with achieving the lowest mean energy overall. This suggests that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, though this is not a strict requirement for the method to work. However, adding a spring constant reintroduces intra-batch dependencies, which means that batch sizes should be of a minimum size for this approach to work effectively. With low batch sizes, this could lead to the corner cutting effects commonly observed in traditional chain-of-states methods, which might make scaling the addition of spring regularization difficult for larger systems. Since we are able to reach the same minimum energy without spring regularization, we recommend against its use, especially when scaling to larger systems where lower batch sizes are preferred.
Spring Constant | MeanMax Energy () | MinMax Energy | Mean Iterations |
-178.22 260.09 | -337.36 | 153533 | |
-155.41 286.63 | -337.36 | 101600 | |
803.61 586.95 | 205.74 | 159033 | |
0.0 | 1460.41 2486.05 | -337.36 | 209133 |
10528.4 7394.89 | 5065.46 | 155400 | |
48639.7 12111.7 | 33143.6 | 107267 |
B.1.4 Network Architecture
We investigated the effect of network depth and width on transition path quality. Table 6 shows that an 8-layer network with a width factor of 4 achieved the best mean energy performance, followed closely by a 6-layer network with width factor 6.
Depth | Width Factor | MeanMax Energy () | MinMax Energy | Mean Iterations |
8 | 4 | 25.02 467.23 | -337.36 | 159333 |
6 | 6 | 72.42 458.72 | -337.36 | 210267 |
8 | 6 | 165.91 1021.79 | -337.36 | 219433 |
8 | 2 | 1434.97 3206.06 | -337.36 | 178167 |
6 | 4 | 1460.34 2486.07 | -337.36 | 209067 |
6 | 2 | 3621.8 3198.25 | 217.83 | 231200 |
4 | 6 | 4242.74 3568.51 | 140.81 | 231967 |
4 | 4 | 5130.75 5758.01 | -328.49 | 228133 |
4 | 2 | 16200.5 8032.18 | 159.21 | 210367 |
B.1.5 Interpolation Type
We examined different interpolation methods for the blending functions that combine the neural network output with the endpoint constraints. Table 7 compares linear, cosine, quintic, and cubic Hermite interpolation. Cosine interpolation performed best in terms of mean energy, though all methods were able to achieve the minimum energy of -337.36 in at least one run, except for quintic. The cubic Hermite interpolation showed higher variance in results.
The specific formulas for each interpolation type are as follows:
Linear interpolation:
Cubic Hermite interpolation:
Quintic interpolation:
Cosine interpolation:
where and are the basis functions that ensure endpoint constraints and , and is the blending function that controls the neural network’s contribution along the path.
Interpolation Type | MeanMax Energy () | MinMax Energy | Mean Iterations |
Cosine | 442.51 769.51 | -337.36 | 198720 |
Linear | 602.61 1057.13 | -337.36 | 219147 |
Quintic | 967.16 1288.57 | -281.14 | 210227 |
Cubic Hermite | 1325.49 2339.75 | -337.36 | 212807 |
B.1.6 AIB9 Network Architecture
Below are the main details of the AIB9 MEP training runs, as shown in the main paper.
Model type: Transition MLP with sinusoidal time embedding
Network depth: 6 hidden layers
Network width: 4 width factor (scaled relative to the number of atoms)
Time embedding: Sinusoidal position encoding with maximum period of 1,000,000
Activation function: ReLU
B.1.7 Training Parameters
Optimizer: Adam
Learning rate:
Batch size: 4
Frames per batch: 32
Spring constant (): 0.0
Interpolation method: Linear
All other parameters are identical to the base settings listed in the ablations section B.1.1.
B.2 BPTI Modeling and Computational Details
B.2.1 System Preparation
As reference conformations, we used the structures from the D.E. Shaw Research BPTI simulation of BPTI (Shaw et al., 2010). The initial and final states for our MEP optimization were selected as the second snapshot (index 1) and the second-to-last snapshot (index -2) from the provided reference structures, as they were the most separated in the CVs determined in the original study, named the backbone RMSD and the disulfide torsion angle.
B.2.2 Explicit Solvent Shell Construction
For the BPTI system, we developed a sophisticated approach to handle explicit water molecules efficiently:
Initial water box generation: Both start and end configurations were solvated with TIP3P water molecules using a padding distance of 2.0 nm and energy minimized for 15,000 steps.
Water shell extraction: Rather than using the entire water box, we extracted spherical shells of water molecules within a 1.4 nm radius of any protein atom for both the start and end configurations. This significantly reduced the system size while maintaining the crucial first hydration shell.
Radius optimization: We employed a binary search algorithm to determine optimal shell radii that would capture an identical number of water molecules in both start and end configurations, ensuring consistent system sizes throughout the transition.
Water molecule matching: We applied the Hungarian algorithm to optimally pair water molecules between the start and end states based on spatial proximity, minimizing the overall displacement of water molecules during the transition, which ensures that the intial guess for the MEP does not have very large energies stemming from overlapping molecules.
B.2.3 Hybrid Neural Network Architecture
For the BPTI system with explicit solvent, we developed a specialized hybrid transition model to handle the larger system size efficiently:
Protein representation: Neural network with 6 layers, a width factor of 4, and sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
Water representation: Parameterized spline interpolation between matched water molecules
Control points: 7 control points for the spline interpolation
Water weight: 0.1 (relative weight of water molecule contributions to the loss)
Clamp threshold: (minimum/maximum energy clipping threshold)
B.2.4 Training Parameters
The BPTI MEP was optimized using the following training parameters:
Optimizer: Adam
Learning rate:
Batch size: 1
Frames per batch: 16
Total iterations: 30,000
All other parameters are identical to the base settings listed in the ablations section B.1.1.
B.2.5 Handling High-Energy Spikes During Training
Physical force fields present unique optimization challenges due to occasional extreme energy spikes when atoms approach one another too closely. To address this, we implemented adaptive gradient clipping in our training procedure:
where is a clipping threshold (set to in our experiments). We specifically use adaptive gradient clipping rather than a fixed threshold because our hybrid architecture contains two fundamentally different parameter types: neural network weights for protein coordinates and spline control points for water molecules. The adaptive nature of the clipping automatically adjusts to the appropriate scale for each parameter type.
Additionally, we apply different scaling factors to these parameter groups:
with for water-related parameters. This differential treatment further balances the influence of each component during optimization. While this approach does not completely eliminate training instabilities, we find it sufficient for most protein systems when combined with careful initialization and learning rate scheduling.
B.2.6 Force Field Parameters
We used the following force field parameters to match the D.E. Shaw simulation conditions:
Protein force field: AMBER99sb-ILDN (amber99sbildn.xml)
Water force field: TIP3P (tip3p.xml)
B.2.7 Computational Performance
For the BPTI system with explicit solvent (approximately 8,300 atoms total):
Training time: Approximately 7 to 15 minutes on a single NVIDIA A6000 GPU
Force field evaluations: Approximately 480,000 evaluations (30,000 iterations x 16 frames per batch)
Trajectory generation: 64 frames for the final MEP visualization and analysis
This computational effort is approximately six orders of magnitude less, in node-hours, than the original D.E. Shaw Research simulation on their specialized Anton supercomputer, which required approximately 1.3 million node-hours to generate the 1.03 millisecond trajectory capturing the same conformational changes.
B.2.8 Computational Resources
All experiments reported in this paper were conducted on a single NVIDIA A6000 GPU with 48GB of memory. Additional computational resources were used during the development phase of this project for hyperparameter tuning across network architectures and optimization parameters.